# Another look at inference after prediction

### v3

@@ Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models increasingly complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data. Motwani and Witten [2023] recently revisited two key PB inference approaches and found that only one method, Prediction-powered Inference (PPI) proposed by Angelopoulos et al. [2023a], achieves (i). In this paper, we find that PPI does not achieve (ii). We revisit the double sampling literature and show that, with a simple modification, PPI can be adjusted to provide theoretically justified improvements in efficiency. We also contextualize PB inference with economics and statistics literature dating back to the 1960s to highlight the utility of classical methods in this contemporary problem. Our extensive theoretical analyses, along with an analysis of UK Biobank data, indicate that our proposal effectively mitigates bias and improves efficiency, making it preferable for use in practice.
@@ 1 Introduction
In the rapidly evolving landscape of data science, accurately predicting an outcome is often just an initial step in scientific inquiry. Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to directly measure, and then drawing statistical inferences from the ML model’s predictions [Hoffman et al., 2024, Fan et al., 2024, Gao et al., 2024, Miao and Lu, 2024]. This prediction-based (PB) inference scenario was introduced by Wang et al in the context of global health and genomics applications, and is comprised of two steps outlined in Figure 1 [Wang et al., 2020]. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw statistical inference on the parameter of interest, such as the association between an outcome and a set of covariates, together with a small amount of gold-standard data  [^1] PB inference has focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et al. [2025]).. In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and ${\bf{X}}$. Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to maximize the use of all available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($N\approx$ 500K individuals), provide an unprecedented data source for high-throughput genetic research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to predict missing outcomes based on fully observed baseline data and the predictions are utilized for genetic discovery [Tong et al., 2020, Alipanahi et al., 2021, Cosentino et al., 2023, Dahl et al., 2023, An et al., 2023, McCaw et al., 2024, Liu et al., 2024, Miao et al., 2024]. Similarly, in the field of structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies [Jumper et al., 2021, Gan et al., 2022]. In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review by experts, enabling downstream epidemiological and clinical studies [Yang et al., 2022, Zhou et al., 2022, Alsentzer et al., 2023].
@@
Given the relevance of PB inference, Motwani and Witten [2023] recently revisited two innovative methods in this field. In the context of least squares estimation, they showed that the approach proposed by Wang et al. [2020] provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios. For example, when the pre-trained model perfectly captures the regression function. This is a significant practical limitation as it is essential that PB inference methods yield valid inference in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) approach proposed by Angelopoulos et al. [2023a] possesses this necessary property and yields valid inference regardless of the quality of the model. This finding is critical given that the PPI paper has been cited over 100 times in its first year of publication in Science.
@@
In this paper, we study the statistical efficiency of the PPI estimator and show that it can be outperformed by traditional inference. We propose a more efficient estimator that can be obtained by simply adding a weight to the PPI estimator, an approach introduced by Chen & Chen (CC) 25 years ago in the context of double sampling designs [Chen and Chen, 2000]. We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten in Section 2. We then study the asymptotic properties of the PPI estimator and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. We also review modern and classical literature from statistics and economics dating back to the 1960s that provides important context for the PB inference problem in Section 4. In Section 5, we evaluate our findings with extensive simulation studies. In Section 6, we study the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of our method. We close with additional discussion in Section 7. All in all, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it the preferred choice for use in PB inference problems.
@@ 2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it on the basis of ${\bf{Z}}$. The ML model, $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$, maps from the predictor space, $\mathcal{Z}$, to the outcome space, $\mathcal{Y}$. The predictions of $Y$ are denoted as $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ be a set of covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. ${\bf{X}}$ may also be predictive of $Y$ and these covariates may or may not be included in ${\bf{Z}}$. We assume that ${\bf{X}}$ and ${\bf{Z}}$ are both fully observed. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
@@
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ are independent and identically distributed samples from the population of interest. More complex scenarios have been the focus of recent work [Angelopoulos et al., 2023a, Zrnic and Candès, 2024a]. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). In fact, $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been proposed by Zrnic and Candès [2024b].
@@ 2.2 Parameter of interest
Following Motwani and Witten [2023], we consider least squares estimation and aim to make inference on the population parameter
@@
which quantifies the association between $Y$ and ${\bf{X}}$. Throughout, we use superscripts to denote different estimators of $\bm{\beta}^{*}$. For example, the simplest unbiased estimator of $\bm{\beta}^{*}$, denoted as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and is the solution to the estimating equation
@@ 2.3 Motivation for PB inference
The goal of PB inference is to utilize the abundant realizations of $\widehat{Y}$, which can be computed in the unlabeled data, to provide valid and efficient inference on the association between $Y$ and ${\bf{X}}$. Motwani and Witten consider the question of validity and we study efficiency in Section 3. With respect to validity, PB inference should enable users to draw correct conclusions regarding the magnitude and direction of $\bm{\beta}^{*}$ and to perform reliable uncertainty quantification. Incorporating information from $\widehat{Y}$ should not change the parameter being estimated and the credibility of inferences should not depend on the quality of predictions from $\hat{f}$.
The need for PB inference is highlighted by limitations of “naive” inference, in which inference on $\widehat{Y}$ is performed in lieu of inference on $Y$ [Wang et al., 2020]. The parameter targeted by naive inference is
$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y}).$
A least squares estimator for $\bm{\gamma}^{*}$, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$, is the solution to
$\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
Critically, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under highly restrictive scenarios, such as when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function. The fundamental problem is that $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ has a different distribution than $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$. To overcome these limitations, PB inference methods utilize $\mathscr{L}$ together with $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
@@ 2.4 PB inference methods
Motwani and Witten revisited two innovative strategies for PB inference. The first is the proposal of Wang et al. [2020], which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches. In their illuminating work, Motwani and Witten show that the estimator of Wang et al is consistent for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. In other scenarios, the Wang et al estimator may not be consistent for $\bm{\beta}^{*}$ even while its standard error converges to 0 as $(n_{\text{\tiny{lab}}},n_{\text{\tiny{unlab}}})\to\infty$. Consequently, the familiar Wald statistic will not converge in distribution under the null, resulting in type I error that is not controlled at the nominal level and confidence intervals with inadequate coverage.
Motwani and Witten also studied the PPI method, which has the crucial property that inference remains valid regardless of the quality of the predictions. Stated otherwise, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency or to yield reliable hypothesis testing and interval estimation. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al. [1994], and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem (e.g., Chen and Chen [2000], Tsiatis [2006], Tian et al. [2012], Wang and Wang [2015]). We return to this point in Section 4.
@@
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right),$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ is the solution to $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ [^2] Angelopoulos et al. [2023a] express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the approach of Wang et al, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. The proposed method therefore asymptotically controls the type I error for Wald-based hypothesis testing and provides confidence intervals with the nominal coverage level.

### v4

@@ Abstract
From structural biology to epidemiology, predictions from machine learning (ML) models are increasingly used to complement costly gold-standard data to enable faster, more affordable, and scalable scientific inquiry. In response, prediction-based (PB) inference has emerged to accommodate statistical analysis using a large volume of predictions together with a small amount of gold-standard data. The goals of PB inference are two-fold: (i) to mitigate bias from errors in predictions and (ii) to improve efficiency relative to traditional inference using only the gold-standard data. While early PB inference methods focused on bias, their ability to enhance efficiency remains unclear. We revisit a popular PB inference method and show that a simple modification can be applied to guarantee improvements in efficiency beyond yielding valid inferences when the ML predictions are imperfect. The utility of this approach in leveraging prediction-based outcomes to enhance efficiency is demonstrated through extensive simulation studies and an application to the UK Biobank data. We further contextualize the problem of PB inference through historical literature from economics and statistics to highlight perspectives from classical methods in this contemporary problem.
@@ 1 Introduction
Researchers are increasingly turning to machine learning (ML) models to obtain predictions for outcomes that are prohibitively time-consuming or expensive to measure directly and conducting statistical analyses treating the predictions as observed data. 1, 2, 3, 4 This prediction-based (PB) inference approach, introduced by Wang et al,5 is comprised of two steps outlined in Figure 1. First, the outcome of interest is imputed using a prediction from a pre-trained model. Then, the predictions are used to draw inferences about the parameters of interest, such as the association parameters between the outcome and a set of covariates, together with a small amount of gold-standard data.***PB inference has primarily focused on outcomes that are difficult to measure. Recent work has considered predicted covariates (e.g., Kluger et al6). In contrast to traditional inference that relies solely on gold-standard data, PB inference aims to leverage the predictions to improve statistical efficiency without introducing bias.
@@
In the last several years, applications of PB inference have arisen in diverse fields, spanning genetics, biology, and medicine. For example, population biobanks, such as the UK Biobank ($n\approx$ 500K individuals), collect high-throughput genomic data on large populations and offer unprecedented opportunities for genetic discovery research. However, many traits and disease outcomes are only partially observed due to the time and cost of measurement. ML models are commonly used to impute missing outcomes based on fully observed baseline data to enable downstream association studies. 7, 8, 9, 10, 11, 12, 13, 14 Similarly, in structural biology, AlphaFold provides structural predictions without the need for expensive and labor-intensive experiments for proteomic studies. 15, 16 In the analysis of electronic health records data, large language models can rapidly ascertain disease status at an accuracy competitive with manual chart review for epidemiological and clinical studies. 17, 18, 19
Figure 1: Prediction-based (PB) inference. $Y$ is the outcome of interest, which is only partially observed, ${\bf{Z}}$ is a set of predictors of $Y$, $\widehat{Y}$ is the prediction of $Y$ based on ${\bf{Z}}$, and ${\bf{X}}$ is a set of covariates. The focus of scientific inquiry is on the association between $Y$ and ${\bf{X}}$. Traditional statistical inference relies solely on the data on $Y$ and ${\bf{X}}$ while PB inference incorporates $\widehat{Y}$ to make use of all of the available data. Dark red represents labeled observations while light red represents unlabeled observations.
@@
Given the relevance of PB inference, Motwani and Witten 20 recently revisited two seminal methods in the field. In the context of least squares estimation, they showed that the post-prediction inference (PostPi) method proposed by Wang et al5 provides a consistent estimator for the association between an outcome and a set of covariates only in restrictive scenarios (e.g., when the pre-trained ML model perfectly captures the underlying regression function). This is a significant practical limitation as it is essential that PB inference methods yield valid inference (i.e., hypothesis tests with controlled type I error and confidence intervals with nominal coverage rates) in the likely scenario that the prediction model is imperfect. In contrast, the prediction powered inference (PPI) method proposed by Angelopoulos et al 21 yields consistent estimation and valid inference regardless of the quality of the ML model. This finding is critical given that the PPI paper, published just over a year ago in the high-impact journal Science, has already been cited nearly 150 times and widely adopted in applications.
In this paper, we build on the work of Motwani and Witten 20 and further examine the PPI estimator by assessing its statistical efficiency. Our contributions are two-fold. First, we develop a set of results that characterize when the PPI estimator gains over traditional estimators using labeled data only. We also show that a more efficient estimator can be obtained by adding a simple weight to the PPI estimator. This approach was proposed by Chen & Chen (CC) over 25 years ago in the context of double sampling designs.22 Second, we revisit literature from statistics and economics and identify work dating back to the 1960s that bears directly on current developments in PB inference. Our review puts classical statistical methods in conversation with this modern problem to inform future methodological advances and applications.
@@
We begin by formally introducing the PB inference problem and reviewing the contributions of Motwani and Witten20 in Section 2. We then study the asymptotic properties of the PPI estimator, showing that it can be outperformed by traditional inference methods, and introduce the CC estimator in Section 3. Our key finding is that the CC estimator is optimal with respect to statistical efficiency in a class of estimators to which the PPI estimator belongs. In Section 4, we discuss related literature and provide connections between the CC estimator and other established methods. We then evaluate our findings with extensive simulation studies in Section 5. In Section 6, we investigate the association between android fat mass and clinical and demographic covariates in the South Asian population in the UK Biobank to demonstrate the practical performance of these approaches. We close with additional discussion in Section 7. Overall, our theoretical and numerical results show that the CC estimator offers a balance between robustness to ML model specification and statistical efficiency, making it a preferred choice for use in practice.
@@ 2.1 Notation and Assumptions
We let $Y$ denote the outcome of interest and ${\bf{Z}}$ a set of predictors. $Y$ is only partially observed and a pre-trained ML model is available to predict it over the full data given ${\bf{Z}}$, which are fully observed. The ML model $\hat{f}:\mathcal{Z}\rightarrow\mathcal{Y}$ maps the predictor space $\mathcal{Z}$ to the outcome space $\mathcal{Y}$ and produces predictions $\widehat{Y}=\hat{f}({\bf{Z}})$. We let ${\bf{X}}=(1,X_{1},\dots,X_{p})^{{\sf\scriptscriptstyle{T}}}$ denote fully observed covariates that are potentially associated with $Y$ and that are the focus of scientific inquiry. ${\bf{X}}$ may also be predictive of $Y$ and these covariates may or may not be included in ${\bf{Z}}$. The available data therefore consists of a labeled dataset and an independent unlabeled dataset, respectively denoted as
@@
We focus on the case in which $\mathscr{L}$ and $\mathscr{U}$ include independent and identically distributed samples from a common population of interest. Additionally, $\hat{f}$ is assumed to be pre-trained with a dataset that is independent of $\mathscr{L}$ and $\mathscr{U}$ and is therefore treated as “fixed” (i.e., the variation in estimating $\hat{f}$ is not of interest in downstream inference). $\hat{f}$ may be a black box function in that its training data and specification are not be available to the users. Approaches that address uncertainty in the estimation of $\hat{f}$ have recently been considered.23
@@ 2.2 Parameter of interest
Following Motwani and Witten,20 we consider least squares estimation and aim to make inferences on the population parameter
@@
which quantifies the association between $Y$ and ${\bf{X}}$. The simplest consistent estimator of $\bm{\beta}^{*}$, which we denote by $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$, uses only $\mathscr{L}$ for estimation and solves the estimating equation
@@ 2.3 Motivation for PB inference
Although $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is consistent and enables valid inference for $\bm{\beta}^{*}$, it has limited precision when gold-standard data is difficult to collect and $n_{\text{\tiny{lab}}}$ is not large. In contrast, the predictions $\widehat{Y}$ can be computed in the full data, leading to a much larger dataset for analysis. However, relying on $\widehat{Y}$ for inference may result in biased estimation depending on the quality of predictions from $\hat{f}$. For example, in “naive” inference, $\widehat{Y}$ is directly substituted for $Y$ in downstream analysis.5 The parameter targeted by this approach is
$\bm{\gamma}^{*}\coloneq\operatorname*{arg\,min}_{\bm{\gamma}}\mathbb{E}[(\widehat{Y}-{\bf{X}}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})^{2}]=[\mathbb{E}({\bf{X}}{\bf{X}}^{{\sf\scriptscriptstyle{T}}})]^{-1}\mathbb{E}({\bf{X}}\widehat{Y})$
and the corresponding least squares estimator, denoted as $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$, is the solution to
$\sum_{i=1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})\coloneq\sum_{i=1}^{n}{\bf{X}}_{i}(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma})=\bm{0}.$
As the distribution of $(\widehat{Y},{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$ may not coincide with that of $(Y,{\bf{X}}^{{\sf\scriptscriptstyle{T}}})^{{\sf\scriptscriptstyle{T}}}$, $\bm{\gamma}^{*}$ bears no necessary correspondence with $\bm{\beta}^{*}$ except under restrictive scenarios (e.g., when ${\bf{Z}}$ contains ${\bf{X}}$ and the ML model captures the true regression function). To overcome these limitations, PB inference methods utilize $\mathscr{L}$ and $\mathscr{U}$ to essentially balance the relative contributions of $Y$ and $\widehat{Y}$ in making inference on $\bm{\beta}^{*}$.
@@ 2.4 PB inference methods
Motwani and Witten20 revisited two seminal strategies for PB inference. The first is the PostPI method of Wang et al,5 which is a two-step procedure to incorporate $\widehat{Y}$ into estimation. First, $\mathscr{L}$ is used to model the relationship between $Y$ and $\widehat{Y}$ to address potential inaccuracies in the predictions. Then, this model is used together with $\mathscr{U}$ to make inference on $\bm{\beta}^{*}$ with either analytical or bootstrap approaches. This approach is shown to yield a consistent estimator for $\bm{\beta}^{*}$ under the strong and unrealistic assumption that the ML model captures the true regression function. When this assumption is not met, the PostPI estimator may not be consistent. Consequently, corresponding hypothesis tests may not have controlled type I error rates and confidence intervals may not have adequate coverage.
Motwani and Witten20 also considered the PPI method,21 which has the crucial property that inference remains valid regardless of the quality of the predictions. That is, the PPI method does not require that $\hat{f}$ capture the true regression function for the estimator to maintain consistency for $\bm{\beta}^{*}$. The proposed estimator utilizes a common statistical technique involving the augmentation of an initial consistent estimator for the parameter of interest with an additional term that converges in probability to 0. This approach dates back to at least Robins et al 24 and has been applied in multiple settings to improve estimation efficiency outside of the PB inference problem. 22, 25, 26, 27 We return to this point in our review of related literature in Section 4.
@@
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}\right)$
where $\hat{\bm{\gamma}}^{\text{\tiny{lab}}}$ solves $\sum_{i=1}^{n_{\text{\tiny{lab}}}}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$ and $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ solves $\sum_{i=n_{\text{\tiny{lab}}}+1}^{n}\bm{\phi}(\widehat{Y}_{i},{\bf{X}}_{i};\bm{\gamma})=\bm{0}$Angelopoulos et al21 express $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ as $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}-\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ where $\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)$ is referred to as a “rectifier.” The rectifier serves the purpose of correcting potential bias in $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ due to inaccuracies in $\widehat{Y}$ for $Y$. We reframe the PPI method to study its statistical efficiency.. In contrast to the PostPI method, Motwani and Witten20 show that $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is consistent for $\bm{\beta}^{*}$ and $n^{\frac{1}{2}}(\hat{\bm{\beta}}^{\text{\tiny{PPI}}}-\bm{\beta}^{*})$ converges weakly to a zero-mean multivariate normal distribution regardless of the quality of the prediction. This approach therefore asymptotically controls the type I error rate in hypothesis tests and produces confidence intervals with the nominal coverage level.